OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (anonymous):

A. an = 896 - 53(n - 2) B. an = 896 - 53(n - 1) C. an = 896 + 53(n + 1) D. an = 896 - 53(n + 1)